OpenStudy (anonymous):

help me with this two geometry questions this is for finding length of a segment (AC) tangent 22/1=14/x .40/1=14/x so far there i want to know how they come up with the .40? . .40(x) =14(1) .40x=14 x=35 and how they come up with the 35 ?

OpenStudy (anonymous):

for the last part x = 14 / 0.40 = 140 / 4 = 35

OpenStudy (anonymous):

have you got the diagram of question?

OpenStudy (anonymous):

diagram drawing you mean?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yes shd i draw it out

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):